The equations represent the demand and supply for silver pendants. The government is considering imposing a $4 per unit tax of buyers of silver pendants.

QD= 50,000-2,000P

QS= -10,000+2,000P

QD (with tax)= 50,000-2,000(P+T)

1. What is the before tax equilibrium price (P) and quantity (Q- in thousands) of pendants?

a. P= $50; Q= 10 thousand

b. P= $20; Q= 15 thousand

c. P= $15; Q= 20 thousand

d. P= $10; Q= 30 thousand.

2. Who carries the highest incidence of taxation in this market?

a. Producers

b. Consumers

c. Government

d. Producers and Consumers have the same tax incidence

3. What is the deadweight loss of this tax?

a. $4000

b. $6000

c. $8000

d. There is not enough information to calculate deadweight loss

4. Suppose the government allows imports into the silver pendants market for the first time.

a. The market price would be higher and domestic production would decrease

b. The market price would be higher and domestic production would increase

c. The market price would be lower and domestic production would decrease

d. The market price would be lower and domestic production would increase.

5. A characteristic of the long run is...

a. there are fixed inputs

b. all inputs can be varied.

c. plant capacity cannot be increased or decreased

d. there are both fixed and variable inputs.
Answer #1

Erice a 18000/ ons of consumer . - surplus ! plus i 15000 loss of producer - Surplus 13000/- 3000 80000 0000 Quantity Тbove 8

1) Before tax equilibrium price is 15000 and quantity is 20000. Option C is correct.

2) Consumers bears the highest burden of tax as loss of consumer surplus rectangle have area 3000 * 16000 = 48000000 while loss of producer surplus have area 2000 * 16000 = 32000000

3) Deadweight loss is the area of triangle of ABC which is 1/2 * 4000 * 5000 = 10000000

4) If imports are allowed inside the country, aggregate demand would fall as AD = C + I + G + Exports - imports.

When AD falls, prices as well as domestic production falls. Thus option C is correct.

5) Option B is correct.